Fluoride ion is added to drinking water to prevent tooth decay. What is the mass (g) of sodium

fluoride in a liter of water that contains 1.193 * 10 19 formula units of the compound? Include a

road map that shows how you planned the solution.


1
Expert's answer
2021-04-21T07:23:26-0400

NaF2+H2O → 2HF + NaOH

1 mole = 6.022 × 1023 formula units

= 1.193 × 1019/6.022 × 1023

= 1.98 × 10-5

= 0.0000198 moles

From the equation, 1mole NaF2 = 1mole H2O

= 0.0000198 mole of H2O

Molar Mass of H2O = 18.01528

= 0.0000198 × 18.01528

= 0.0003567g of NaF